OpenStudy (anonymous):

find the difference of 9srqt7 and sqrt28

OpenStudy (anonymous):

\[7\sqrt{7}\]

OpenStudy (anonymous):

\[9\sqrt{7}-\sqrt{28}=9\sqrt{7}-2\sqrt{7}=7\sqrt{7}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

answer is \[7\sqrt{7}\]